3. Lucas is the worlds WORST sno cone maker. He drops 78% of the snow cones he makes. Its a terrible waste of sno. Anyway,
Answer #1

Let~X=no.~of~dropped~Sno~cones.\\ Then~X\sim Binomial(20,0.78).\\\\ (a)~P(X=12)=0.0351\\\\ (R~code:~round(dbinom(12,20,0.78),4)).\\\\ (b)~P(X=20)=0.0069\\\\ (R~code:~round(dbinom(20,20,0.78),4)).\\\\ (c)~P(X<8)=P(X\leq 7)=0.0000\\\\ (R~code:~round(pbinom(7,20,0.78),4)).\\\\ (d)~P(X\geq 15)=1-P(X\leq 14)=0.7343\\\\ (R~code:~round(1-pbinom(14,20,0.78),4)).\\\\